Assuming the chance of each hit is independent, why do you think they wouldn't be additive?
P(hits >= 1) = 1 - P(hits = 0).
Solving for the left side of the equation directly gets trickier: P(hits >= 1) = P(hits = 1) + P(hits = 2) + ... + P(hits = k)
P(hits = 1) = 100 chose 1 * 0.01^1 * 0.99^99
...
p(hits = n) = 100 chose n * 0.01^n *0.99^(100-n)
Taking that sum where n=100 yields the same as 1 - P(hits = 0)
= 1 - (100 chose 0 *0.1^0 *0.99^100)
= 1 - 0.99^100
[0] https://en.wikipedia.org/wiki/Binomial_distributionThe easiest way to solve "at least once"-type questions is invariably by subtracting P(!"at least once") = P("zero hits") from 1.
To do it with addition, you'd have to list all possible outcomes completely. Those would, individually, be far lower probabilities. Example:
P("at least one "six" when rolling two dice")
= P("first not six, second six") + P("first six, second not six") + P("first six, second also six")
= (1/6 * 5/6) + (5/6 * 1/6) + (1/6 * 1/6)
= 11/36
Shorter: 1 - P("never six")
= 1 - P("first not six, second not six")
= 1 - (5/6 * 5/6) = 36/36 - 25/36
= 11/36
The number of possible outcomes you have to list and add together is 2^<number of trials> - 1. For two dices, that's a manageable 2^2 - 1 = 1. For 100 rockets it's 2^100 - 1 = 1.2676506E30 (- 1).Think of it like flipping coins. For a given coin flip, 50% odds of heads and 50% odds of tails.
But if you flip the coin 100000000 times there's still a tiny outside chance of tails coming up every time.
As you flip more coins, the odds of getting a heads on at least one of them will asymptotically approach 100%, but it will never get there.